I'm having trouble with a math problem about systems of linear equations. FInd out the values of x and y and please show your work.
7x+3y=1
4x-y=6
I thank you for the help.

Answer by stanbon(75887)

7x+3y=1
4x-y=6
----
Multiply thru the bottom equation by 3:
-------
7x + 3y = 1
12x -3y = 18
---------------------
Add and solve for "x":
19x = 19
x = 1
----
Solve for "y":
4x-y = 6
4 - y = 6
y = -2
==================
Cheers,
Stan H.
===================